If you increase the dimensions of a square pyramid by 6 times, how much larger will the new surface area be?
1
Expert's answer
2017-02-22T10:01:05-0500
S=S_base+S_side
S_base=a^2
S_side=4(1/2 a×h)
h=√(a^2-a^2/4)
S=36+36√3=36(1+√3)